How many Tm2 are there in 4.14 x 1025 yd2?


1
Expert's answer
2021-02-11T04:54:40-0500

1 yd2 = 8.361×10-25 Tm2

4.14×1025 yd2 = 4.4×1025 × 8.361×10-25 Tm2

= 36.7884 Tm2